OpenStudy (anonymous):

multiply (x-7/3)(x+3/7) simplify

OpenStudy (anonymous):

x^2-(40 x)/21-1

OpenStudy (anonymous):

\[(x-\frac{7}{3})(x+\frac{3}{7})=x^2-\frac{7}{3}x+\frac{3}{7}x-\frac{7}{3}\times \frac{3}{7}\]

OpenStudy (anonymous):

multiply, x by x x by 3/7 -7/3 by x and -7/3 by 3/7

OpenStudy (anonymous):

= what mathhelp said

OpenStudy (saifoo.khan):

hey, there?

OpenStudy (saifoo.khan):

@Rachel, can u please tell me how to put these in compund form?

OpenStudy (saifoo.khan):

4>-2x + 3 1/2 > -x x < 1/2 ======================= 9=<-3x + 4 4/3 <=-x x => -4/3

OpenStudy (anonymous):

\[4>-2x+3\] \[2x>-1\] \[x>-\frac{1}{2}\] and \[ 9\geq-3x+4\] \[3x\geq -5\] \[x\geq-\frac{5}{3}\] so i guess if both conditions must be satisfied the answer is \[x>-\frac{1}{2}\]

OpenStudy (saifoo.khan):

HELLO?

OpenStudy (anonymous):

hello?